OpenStudy (anonymous):

Can someone show me how to do this problem? Mom folded the napkins into fourths, so they would be 8 inch squares. What is the area of the napkins when they are fully open?

OpenStudy (anonymous):

A = 8*8 = 64 64 * 4 squares = 256 inches squared